I have my exams in 12 days.. Please help me to complete chemistry... By telling some important concepts only... 

Grade:12th pass

3 Answers

sree lakshmi
3 years ago
students must cover all the syllabus to score good marks in exam Since questions in exam are asked from all the topics but still the most important topics are

Atomic Structure

Gaseous & Liquid State

Electrochemistry

Chemical Kinetics

Nuclear & Surface Chemistry

Chemical Bonding

d-Block Elements

Co-ordination Compound

Alkyl Halides

Aromatic Compounds

Thermodynamics

p-Block Elements

General Organic Chemistry

Isomerism

Aldehydes and Ketones

Carboxylic Acids & Its Derivatives

if you dont know the concepts in organic chemistry atleast try to memmorise organic name reactions 

Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

If the guard at the rear of B is passing the driver at the front of A then B has not only caught up the initial distance between the two trains but has also completely overtaken train A.
Let the original separation distance be x. Then distance to be covered starting with relative velocity zero is s = 800 + x (800m being the overtaking distance from front meets rear to rear leaves front).
Distance s covered at acceleration a starting from (relative) speed zero = 1/2 at^2 which for t=50 gives:
= 0.5 x 50^2 = 1250m
So 800 + x = 1250 Therefore x = 1250 -800 = 450m.

Thanks and Regards
adilakshmigogada 27
3 years ago

Atomic Structure

Gaseous & Liquid State

Electrochemistry

Chemical Kinetics

Nuclear & Surface Chemistry

Chemical Bonding

d-Block Elements

Co-ordination Compound

Alkyl Halides

Aromatic Compounds

Thermodynamics

p-Block Elements

General Organic Chemistry

Isomerism

Aldehydes and Ketones

Carboxylic Acids & Its Derivatives